Santan bushes are to be planted. around the circumference of the circular garden in radius of 5 m at 3 dm apart. how many bushes will be planted?
Answer by sachi(548)

radius= 5 m
so circumference of the circular garden=2*pi*r=2*3.14*5=31.4m
Santan bushes are to be planted at 3 dm apart i.e 0.3m apart
so no of bushes required=31.4/0.3=104 approx
ans